Shift summary, Thursday Oct. 28 DAY shift

Vivian O'Dell ((no email))
Thu, 28 Oct 1999 16:21:40 -0500

Shift Summary for Day shift, 10/28/99

On Shift: Vivian O'Dell (walk through)

A variety of other committed collaborators were working throughout this
shift.

Hours of beam delivered at all intensities: 0.0
Hours of beam delivered at nominal intensity: 0.0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

No beam this shift due to TeV studies.

Activities:
08:00 Walkthrough. Nothing obvious wrong.

12:30 Alcohol bubbler filled. Was 2.25" now 3.5".

13:00 Ed, Doug, Sasha and some other guy I didn't know
made an access of the vacuum tank and found that the
downstream west bearing had disintegrated.

13:30 Walkthrough. Nothing obvious wrong. Did the gas check
list. DN was playing with the HE bags so the pressure was
nearly 0... Don't panic!

14:35 2 new bottles of AE now online west side.

15:00 Duane from SGI installs R10k loaner board on ktev4.

that's it...

=========================================================================
Runs ended this shift:

Tapes brought to FCC this shift: 0